# Medical Image Classification with KAN-Integrated Transformers and Dilated Neighborhood Attention

### v1

@@ Abstract
Convolutional networks, transformers, hybrid models, and Mamba-based architectures have demonstrated strong performance across various medical image classification tasks. However, these methods were primarily designed to classify clean images using labeled data. In contrast, real-world clinical data often involve image corruptions that are unique to multi-center studies and stem from variations in imaging equipment across manufacturers. In this paper, we introduce the Medical Vision Transformer (MedViTV2), a novel architecture incorporating Kolmogorov-Arnold Network (KAN) layers into the transformer architecture for the first time, aiming for generalized medical image classification. We have developed an efficient KAN block to reduce computational load while enhancing the accuracy of the original MedViT. Additionally, to counteract the fragility of our MedViT when scaled up, we propose an enhanced Dilated Neighborhood Attention (DiNA), an adaptation of the efficient fused dot-product attention kernel capable of capturing global context and expanding receptive fields to scale the model effectively and addressing feature collapse issues. Moreover, a hierarchical hybrid strategy is introduced to stack our Local Feature Perception and Global Feature Perception blocks in an efficient manner, which balances local and global feature perceptions to boost performance. Extensive experiments on 17 medical image classification datasets and 12 corrupted medical image datasets demonstrate that MedViTV2 achieved state-of-the-art results in 27 out of 29 experiments with reduced computational complexity. MedViTV2 is 44% more computationally efficient than the previous version and significantly enhances accuracy, achieving improvements of 4.6% on MedMNIST, 5.8% on NonMNIST, and 13.4% on the MedMNIST-C benchmark. Our code is available at https://github.com/Omid-Nejati/MedViTV2.git
@@
Medical image classification , Kolmogorov–Arnold Networks , Medical image corruption , Deep learningbooktabs
1 Introduction
Computer-aided diagnosis (CAD) systems have attracted significant research interest in medical image analysis, aiming to assist clinicians in making diagnostic decisions. These systems are applied to various modalities, including X-ray radiography [1], computed tomography (CT) [2], magnetic resonance imaging (MRI) [3], ultrasound [4], and digital pathology [5]. The success of deep learning in this domain is partly attributed to the increasing availability of large-scale datasets. Large datasets with reliable labels are ideal for training deep neural networks. However, collecting labeled medical images remains challenging due to data privacy concerns and the time-consuming nature of expert annotations.
@@ 1 Introduction
While MedViTV1 demonstrated strong performance on medical image classification benchmarks, it exhibits weaknesses in model scalability, which is instrumental for more complex data and tasks, with accuracy dropping as the number of parameters increases, as shown in Figure 1. This motivates us to explore components that can support large-scale training, boosting the expressiveness and efficiency of our model while enhancing its competitiveness across a wide range of medical benchmarks. Recently, Dilated Neighborhood Attention (DiNA) [21] introduced an efficient and scalable sliding window attention mechanism in vision tasks. DiNA is a pixel-wise operation that localizes self-attention to the nearest neighboring pixels, enabling linear time and space complexity. The sliding window pattern allows DiNA to capture more global context and exponentially expand receptive fields at no additional cost. Additionally, Kolmogorov-Arnold Networks (KANs) [22] have emerged as a powerful alternative to multi-layer perceptrons (MLPs). In most MLP-based neural networks, function combinations occur within the activation functions. In contrast, KANs perform these combinations directly on the functions that map inputs to outputs. A few recent studies have begun exploring the effectiveness of incorporating KAN layers into transformers, demonstrating that this can boost the expressiveness and efficiency of transformers [23, 24], thereby enhancing their competitiveness across a wide range of applications.
@@
To the best of our knowledge, this is the first study that integrates the KAN into the feed-forward pathway of transformers for medical image classification. This adaptation reduces computational complexity by 44% while significantly enhancing performance.
@@ 2 Related Works
Medical image classification remains a significant challenge, critical in organizing large volumes of data into meaningful categories [25]. Over the past decade, CNNs have dominated the field of image classification. They have been widely employed in applications such as cancer detection [26], skin disease diagnosis [27], thoracic surgery [28], retinal disease identification [29], and fetal brain volume estimation [30].
More recently, Vision Transformers (ViTs) have emerged as a powerful alternative to conventional CNNs, achieving remarkable success in various image classification tasks [31, 32, 33, 34]. ViTs offer several advantages, including the ability to model long-range dependencies, adapt to diverse inputs, and generate attention maps that highlight critical regions within an image [35]. These features have sparked significant interest in leveraging Transformer-based models for medical image classification, where precise classification is increasingly essential to support timely clinical decision-making, particularly for difficult cases. Early ViT models typically rely on large-scale datasets and relatively simple configurations [36]. However, recent advancements have integrated inductive biases related to visual perception into ViT architectures [37, 38, 39, 40]. This evolution has made ViTs more adaptable and effective in classification, registration, and segmentation. By treating images as sequences of patches without incorporating 2D inductive biases, ViTs are particularly suitable for multimodal applications [41]. In particular, the growth of datasets and innovations in model architectures have driven ViT-based foundation models with unprecedented capabilities, enabling flexible applications in medical imaging [42, 43, 44]. For example, researchers have introduced FastGlioma, a tool for detecting tumor infiltration during surgery [45], while RETFound learns generalizable representations from unlabelled retinal images, enabling label-efficient model adaptation across various applications [46]. Additional applications include leveraging tumor registry data and demographic information to predict overall survival rates [47].
@@ 2.2 Kolmogorov–Arnold Networks
KANs have inspired numerous studies that demonstrate their effectiveness across various domains, including keyword spotting, complex optimization problems, survival analysis, time series forecasting, quantum computing, and vision tasks [48, 49, 50]. Furthermore, many advanced KAN models leverage well-established mathematical functions, particularly those adept at handling curves, such as B-splines [51], which combine B-splines and Radial Basis Functions (RBF) to fit input data during training. FastKAN [52] approximates third-order B-splines in KANs using Gaussian RBF, while DeepOKAN [53] directly employs Gaussian RBF instead of B-splines. Other approaches, such as FourierKAN-GCF [54], wavelet-based KANs [55], and polynomial-function-based KANs [56], explore the suitability of various basis functions in KAN models for classification and other tasks.
@@ 3.2 Dilated Attention Block
To introduce an efficient attention into LFP block of our model at no additional cost, we use dilated attention as shown in Figure 2. Dilated attention confines self-attention to the nearest neighbors of each pixel, maintaining the same network complexity as well as parameter count as shifted windows attention [59]. However, it operates within overlapping shifted windows, thereby preserving translation equivariance. Specifically, let denote a scaling parameter and represent the embedding dimension. In DiNA, for a given dilation value , we determine as the nearest neighbor of token that satisfies the condition . Using this definition, the -dilated neighborhood attention weights for the token, with a neighborhood size , , can be expressed as follows:
@@ 3.3 Kolmogorov–Arnold Networks (KANs)
C C C C | 31.1 | 7.2 | 23.3 | 68.76 |
C C C T | 32.4 | 7.4 | 23.5 | 68.93 |
C C T T | 34.5 | 7.8 | 24.6 | 68.37 |
C T T T | 34.8 | 7.8 | 24.9 | 68.28 |
C C C H | 30.9 | 6.9 | 18.8 | 68.80 |
C C H H | 32.5 | 7.4 | 19.4 | 69.58 |
C H H H | 33.1 | 7.6 | 19.6 | 69.73 |
C H H T | 33.4 | 7.6 | 20.1 | 70.51 |
@@ 4.7 Heatmap Visualization
To gain deeper insight into the learning behavior, we perform a qualitative analysis of the feature space, as shown in Figure 4. Using both MedViTV1 and MedViTV2 in their tiny and large configurations, we generate heatmaps for several datasets from the MedMNIST benchmark using Grad-CAM [93]. An intriguing phenomenon, referred to as ”feature collapse” [94], is observed in MedViTV1-L for certain datasets, including BreastMNIST, RetinaMNIST, and TissueMNIST. This occurs when many feature maps become saturated or inactive, primarily in the dimension expansion layers of the MedViTV1 block. To address this issue, we propose combining new blocks to diversify the feature representations during scaling, effectively mitigating feature collapse. As a result, MedViTV2-L demonstrates improved attention quality, focusing on more relevant areas of the images compared to its smaller version. Notably, MedViTV2-T has only 12 million parameters and captures critical features, highlighting the most important regions. In contrast, MedViTV2-L, with over ten times the number of parameters, is capable of capturing richer features, resembling segmentation masks in cases such as RetinaMNIST and BreastMNIST.

### v2

# MedViT V2: Medical Image Classification with KAN-Integrated Transformers and Dilated Neighborhood Attention

@@ Abstract
Convolutional networks, Transformers, hybrid models, and Mamba-based architectures have demonstrated strong performance across various medical image classification tasks. However, these methods were primarily designed to classify clean images using labeled data. In contrast, real-world clinical data often involve image corruptions that are unique to multi-center studies and stem from variations in imaging equipment across manufacturers. In this paper, we introduce the Medical Vision Transformer (MedViTV2), a novel architecture incorporating Kolmogorov-Arnold Network (KAN) layers into the Transformer architecture for the first time, aiming for generalized medical image classification. We have developed an efficient KAN block to reduce computational load while enhancing the accuracy of the original MedViT. Additionally, to counteract the fragility of our MedViT when scaled up, we propose an enhanced Dilated Neighborhood Attention (DiNA), an adaptation of the efficient fused dot-product attention kernel capable of capturing global context and expanding receptive fields to scale the model effectively and addressing feature collapse issues. Moreover, a hierarchical hybrid strategy is introduced to stack our Local Feature Perception and Global Feature Perception blocks in an efficient manner, which balances local and global feature perceptions to boost performance. Extensive experiments on 17 medical image classification datasets and 12 corrupted medical image datasets demonstrate that MedViTV2 achieved state-of-the-art results in 27 out of 29 experiments with reduced computational complexity. MedViTV2 is 44% more computationally efficient than the previous version and significantly enhances accuracy, achieving improvements of 4.6% on MedMNIST, 5.8% on NonMNIST, and 13.4% on the MedMNIST-C benchmark.
@@
Medical image classification , Kolmogorov–Arnold Networks , Medical image corruption , Deep learning1 Introduction
Computer-aided diagnosis (CAD) systems have attracted significant research interest in medical image analysis, aiming to assist clinicians in making diagnostic decisions. These systems are applied to various modalities, including X-ray radiography [1], computed tomography (CT) [2], magnetic resonance imaging (MRI) [3], ultrasound [4], and digital pathology [5]. The success of deep learning in this domain is partly attributed to the increasing availability of large-scale datasets. Large datasets with reliable labels are ideal for training deep neural networks. Although, collecting labeled medical images remains challenging due to data privacy concerns and the time-consuming nature of expert annotations.
@@
While MedViTV1 demonstrated strong performance on medical image classification benchmarks, it exhibits weaknesses in model scalability, which is instrumental for more complex data and tasks, with accuracy dropping as the number of parameters increases, as shown in Figure 1. This motivates us to explore components that can support large-scale training, boosting the expressiveness and efficiency of our model while enhancing its competitiveness across a wide range of medical benchmarks. Recently, Dilated Neighborhood Attention (DiNA) [21] introduced an efficient and scalable sliding window attention mechanism in vision tasks. DiNA is a pixel-wise operation that localizes self-attention to the nearest neighboring pixels, enabling linear time and space complexity. The sliding window pattern allows DiNA to capture more global context and exponentially expand receptive fields at no additional cost. Additionally, Kolmogorov-Arnold Networks (KANs) [22] have emerged as a powerful alternative to multi-layer perceptrons (MLPs). In most MLP-based neural networks, functional combinations occur within the activation functions. In contrast, KANs directly perform these combinations on the functions that map inputs to outputs. As a result, KANs rely on simpler functional modules rather than complex, heavily parameterized architectures. This reduction in complexity enables training on medical datasets, which are typically smaller than computer vision datasets. Recent studies have begun to explore the effectiveness of incorporating KAN layers into Transformers, demonstrating that this integration can enhance both the expressiveness and efficiency of Transformers [23, 24], thereby improving their competitiveness across a wide range of applications. Building on this, we incorporate KAN layers into MedViT with the goal of effectively capturing complex non-linear medical structures, while their functional decomposition helps separate and capture the true signal from the noise, which is often prevalent in medical images.
@@
To the best of our knowledge, this is the first study that integrates KAN into the feed-forward pathway of Transformers for medical image classification. This adaptation reduces computational complexity by 44% while significantly enhancing performance.
@@
Our code is available at https://github.com/Omid-Nejati/MedViTV2.git
@@ 2 Related Works
Medical image classification remains a significant challenge, critical in organizing large volumes of data into meaningful categories [25]. Over the past decade, CNNs have dominated the field of image classification. They have been widely employed in applications such as cancer detection [26], skin disease diagnosis [27, 28], thoracic surgery [29], retinal disease identification [30], and fetal brain volume estimation [31].
More recently, Vision Transformers (ViTs) have emerged as a powerful alternative to conventional CNNs, achieving remarkable success in various image classification tasks [32, 33, 34, 35]. ViTs offer several advantages, including the ability to model long-range dependencies, adapt to diverse inputs, and generate attention maps that highlight critical regions within an image [36]. These features have sparked significant interest in leveraging Transformer-based models for medical image classification, where precise classification is increasingly essential to support timely clinical decision-making, particularly for difficult cases. Early ViT models typically rely on large-scale datasets and relatively simple configurations [37]. However, recent advancements have integrated inductive biases related to visual perception into ViT architectures [38, 39, 40, 41]. This evolution has made ViTs more adaptable and effective in classification, registration, and segmentation. By treating images as sequences of patches without incorporating 2D inductive biases, ViTs are particularly suitable for multimodal applications [42, 43]. In particular, the growth of datasets and innovations in model architectures have driven ViT-based foundation models with unprecedented capabilities, enabling flexible applications in medical imaging [44, 45, 46]. For example, researchers have introduced FastGlioma, a tool for detecting tumor infiltration during surgery [47], while RETFound learns generalizable representations from unlabelled retinal images, enabling label-efficient model adaptation across various applications [48]. Additional applications include leveraging tumor registry data and demographic information to predict overall survival rates [49].
@@ 2.2 Kolmogorov–Arnold Networks
KANs have inspired numerous studies that demonstrate their effectiveness across various domains, including keyword spotting, complex optimization problems, survival analysis, time series forecasting, quantum computing, and vision tasks [50, 51, 52]. Furthermore, many advanced KAN models leverage well-established mathematical functions, particularly those adept at handling curves, such as B-splines [53], which combine B-splines and Radial Basis Functions (RBF) to fit input data during training. FastKAN [54] approximates third-order B-splines in KANs using Gaussian RBF, while DeepOKAN [abueidda2024deepokan] directly employs Gaussian RBF instead of B-splines. Other approaches, such as FourierKAN-GCF [xu2024fourierkan], wavelet-based KANs [seydi2024unveiling], and polynomial-function-based KANs [58], explore the suitability of various basis functions in KAN models for classification and other tasks.
@@ 3.2 Dilated Attention Block
To introduce an efficient attention into LFP block of our model at no additional cost, we use dilated attention as shown in Figure 2. Dilated attention confines self-attention to the nearest neighbors of each pixel, maintaining the same network complexity as well as parameter count as shifted windows attention [32]. Nonetheless, it operates within overlapping shifted windows, thereby preserving translation equivariance. Specifically, let denote a scaling parameter and represent the embedding dimension. In DiNA, for a given dilation value , we determine as the nearest neighbor of token that satisfies the condition . Using this definition, the -dilated neighborhood attention weights for the token, with a neighborhood size , , can be expressed as follows:
@@ 3.3 Kolmogorov–Arnold Networks (KANs)
C C C C | 27.4 | 7.2 | 23.3 | 68.76 |
C C C T | 28.5 | 7.4 | 23.5 | 68.93 |
C C T T | 30.7 | 7.8 | 24.6 | 68.37 |
C T T T | 31.0 | 7.8 | 24.9 | 68.28 |
C C C H | 27.1 | 6.9 | 18.8 | 68.80 |
C C H H | 28.6 | 7.4 | 19.4 | 69.58 |
C H H H | 29.3 | 7.6 | 19.6 | 69.73 |
C H H T | 29.6 | 7.6 | 20.1 | 70.51 |
Hybrid Strategy | Model | Configs | Complexity | Latency(ms) | Acc(%) | ||||
S1 | S2 | S3 | S4 | param(M) | FLOPs(G) | ||||
Small | 2 | 2 | (8+1) | 2 | 27.0 | 6.9 | 20.4 | 68.35 | |
2 | 2 | (11+1) | 2 | 29.1 | 7.7 | 20.9 | 69.59 | ||
2 | 2 | (14+1) | 2 | 31.1 | 8.5 | 21.6 | 69.80 | ||
Small | 2 | 2 | (1+1)3 | 2 | 31.5 | 8.3 | 20.7 | 69.31 | |
2 | 2 | (2+1)3 | 2 | 33.5 | 9.1 | 21.2 | 70.54 | ||
2 | 2 | (3+1)3 | 2 | 35.5 | 9.8 | 21.5 | 70.54 | ||
2 | 2 | (2+1)2 | 2 | 29.6 | 7.6 | 20.1 | 70.51 | ||
2 | 2 | (3+1)2 | 2 | 30.9 | 8.2 | 20.7 | 70.52 |
@@ 3.4 Hierarchical Hybrid Strategy
To further assess the generalized efficacy of , we scaled up a MedViT model by augmenting the block count in its third stage. The initial three rows of experimental results in Table 2 reveal that a mere increase in the number of blocks yields diminishing performance returns, eventually leading to saturation. This observation implies that a straightforward expansion of model size by simply increasing the value of in the pattern, i.e., adding more convolution blocks, is not the most effective approach. Prompted by this, we conducted extensive experiments to meticulously investigate the impact of on model performance. As presented in the middle section of Table 2, we designed models with diverse configurations for the third stage. To facilitate a fair comparison based on similar latency profiles, we employed stacked groups of the pattern when was small. Remarkably, we discovered that models employing the stacked pattern achieved superior performance compared to those using only a single pattern. This suggests that the iterative and principled combination of low-frequency signal extractors (LFP) and high-frequency signal extractors (GFP) through the sub-pattern is crucial for learning higher-quality representations. From Table 2, the model with in the third stage demonstrates the optimal balance between performance and latency. Subsequently, we constructed even larger models by increasing the value of within the pattern for the third stage. As indicated in Table 2, for small size of our model, setting and yielded the best results. However, the performance improvement over and was marginal, while simultaneously introducing a significant increase in complexity. This observation, combined with the broader analysis, confirms the wide applicability and effectiveness of our proposed pattern. Consequently, we adopted as the default configuration for the remainder of this paper due to its favorable performance-complexity trade-off.
Based on the aforementioned Hierarchical Hybrid Strategy, we constructed MedViT, formally defined as:
where denotes the stage index. refers to the LFP at stage . represents an identity layer for the first stage () and the GFP for subsequent stages (). The operator signifies the sequential concatenation of these stages.
@@ 4.7 Heatmap Visualization
To gain deeper insight into the learning behavior, we perform a qualitative analysis of the feature space, as shown in Figure 4. Using both MedViTV1 and MedViTV2 in their tiny and large configurations, we generate heatmaps for several datasets from the MedMNIST benchmark using Grad-CAM [93]. Also, we perform qualitative analysis in the feature space. We visualize the activations in Figure 5. An intriguing phenomenon, referred to as “feature collapse” [94], is observed in MedViTV1-L for certain datasets, including BreastMNIST, RetinaMNIST, and TissueMNIST. This occurs when many feature maps become saturated or inactive, primarily in the dimension expansion layers of the MedViTV1 block. To address this issue, we propose combining DiNA block to diversify the feature representations during scaling, effectively mitigating feature collapse. As a result, MedViTV2-L demonstrates improved attention quality, focusing on more relevant areas of the images compared to its smaller version. Notably, MedViTV2-T has only 12 million parameters and captures critical features, highlighting the most important regions. In contrast, MedViTV2-L, with over ten times the number of parameters, is capable of capturing richer features, resembling segmentation masks in cases such as RetinaMNIST and BreastMNIST.
@@ 5.1 Impact of Components
Model | Memory usage | FLOPs | Receptive Field | TissueMNIST(%) | TissueMNIST-C(%) |
ViT | 64.7 | 46.3 | |||
(6.2 GB) | (17.6 G) | ||||
Swin | 66.3 | 39.7 | |||
(6.7 GB) | (15.4 G) | ||||
NA | 69.5 | 49.1 | |||
(5.0 GB) | (13.7 G) | ||||
DiNA | 71.1 | 54.5 | |||
(5.2 GB) | (15.6 G) |
5.2 DiNa vs. Attention Methods
To fully appreciate DiNA’s effectiveness, particularly in comparison to other models, a thorough analysis of its receptive field (RF) is essential. Table 11 provides a comparative overview of receptive field sizes across various attention patterns, along with their associated computational (FLOPs) and memory costs for TissueMNIST and TissueMNIST-C. We determine the receptive field size by considering the number of layers , kernel size , and the number of tokens . Initially, Neighborhood Attention (NA) has a receptive field of , which expands by with each subsequent layer while maintaining the central pixel fixed. Conversely, the Window Self Attention employed in the Swin Transformer [32] intrinsically maintains a constant receptive field because its window partitioning prevents interactions between windows, thereby precluding further RF expansion. This limitation is mitigated by pixel-shifted WSA, which facilitates a receptive field expansion of exactly one window per layer, translating to a expansion per layer.
In stark contrast to the fixed receptive field growth exhibited by NAT and Swin, DiNA’s receptive field demonstrates remarkable adaptability and is contingent on the chosen dilation. Its receptive field can span from a linear progression, similar to NAT’s (when all dilation values are set to 1), to an exponential growth of (when dilation increases progressively). This inherent flexibility in dilation is a primary driver of DiNA’s superior performance. Regardless of the dilation strategy, the first layer of DiNA invariably yields a receptive field of size . When sufficiently large dilation values are utilized, each DiNA layer contributes a receptive field of size per dimension, culminating in an effective receptive field of . This allows for DiNA and NA combinations, when paired with optimal dilation values, to achieve an impressive exponential receptive field expansion up to . This observation is consistent with established knowledge regarding dilated convolutions, which are known to exhibit exponential receptive field growth when using exponentially increasing dilation factors [21]. In our experiments, DiNA demonstrates notable improvements, surpassing Swin and NAT by 4.8% and 1.6%, respectively, on TissueMNIST. Furthermore, DiNA significantly excels on the corruption benchmark, outperforming Swin and NAT by 14.8% and 5.1%, respectively, on TissueMNIST-C.
Dilation per level | TissueMNIST(%) | DermaMNIST(%) | TissueMNIST-C(%) | DermaMNIST-C(%) |
1, 1, 1, 1 | 68.7 | 78.8 | 45.6 | 52.0 |
2, 2, 2, 1 | 69.8 | 79.2 | 49.4 | 60.9 |
4, 4, 2, 1 | 70.4 | 80.1 | 52.3 | 65.5 |
8, 4, 2, 1 | 71.1 | 80.8 | 54.5 | 69.6 |
5.3 Dilation Values
Our findings in Table 12 illustrate the impact of different dilation values on both standard and robust classification performance. It should be highlighted that the progressively increasing dilation sequence was exclusively utilized for the classification task. This is because, from a theoretical standpoint, the input feature maps must have dimensions at least equal to the product of the kernel size and the dilation factor. Therefore, for MedMNIST images at a resolution, the sequence represents the upper limit of feasible dilation. Should image resolution be increased, higher dilation values could be employed. We investigated a spectrum of dilation configurations within DiNA layers, ranging from the NAT model’s uniform to MedViTV2’s varied , as summarized in Table 12. Ultimately, we opted for a “gradual” dilation scheme, where the dilation value incrementally approaches a predetermined maximum.
Model | #Param. | FLOPs | Memory usage | throughput(ms) | TissueMNIST(%) | TissueMNIST-C(%) | |
MLP | 78.2 M | 7.5 G | 3.7 GB | 178 | 65.9 | 43.2 | |
EfficientKAN | 90.1 M | 16.8 G | 17.8 GB | 521 | 69.5 | 45.9 | |
FastKAN | 84.3 M | 16.5 G | 7.6 GB | 309 | 69.7 | 43.4 | |
FC-KAN | 97.6 M | 18.3 G | 19.5 GB | 576 | 70.8 | 46.7 | |
Our | 72.3 M | 15.6 G | 5.2 GB | 293 | 71.1 | 54.5 |
5.4 Impact of KAN Block
To validate the effectiveness of our proposed KAN component, we substituted the KAN component in MedViTV2 with several established blocks, including the MLP from ViT [37], EfficientKAN222https://github.com/Blealtan/efficient-kan, FastKAN [54], and FC-KAN [95]. For a rigorous and fair comparison, we consistently applied LFP and GFP in the construction of all models, adhering to the settings outlined in Section 4.2. As demonstrated in Table 13, our proposed KAN achieves the most favorable balance between resource consumption and accuracy across both benchmarks. This empirically confirms the advantages of our KAN component. For example, our KAN outperforms the recent FC-KAN block [95] by 0.3% on TissueMNIST and a substantial 7.8% on TissueMNIST-C, while concurrently achieving a significant reduction in memory usage by approximately 73.33%.
5.5 Feature Collapse Analysis
Several recent works have mathematically analyzed the feature representation capabilities of deep models [96, 97, 94]. To quantitatively validate our observations, and building upon prior research, we conducted a feature cosine distance analysis. Given an activation tensor , where denotes the feature map of the -th channel, we reshape each feature map into a -dimensional vector. We then compute the average pairwise cosine distance across all channels using the formula: . In this context, a higher distance value signifies greater feature diversity, while a lower value indicates feature redundancy.
For this analysis, we randomly selected 5,000 images from various classes within the PathMNIST validation set. High-dimensional features were extracted from each layer of different models, including the pre-trained Swin model [32], the pre-trained ViT model [37], the pre-trained NAT model [21], and the pre-trained MedViTV1 [19]. Subsequently, we computed the cosine distance per layer for each image and averaged these values across all images. The aggregated results are presented in Figure 6. Consistent with our observations from earlier feature map visualizations, the pre-trained MedViTV1 model clearly exhibits a tendency towards feature collapse. This finding strongly motivates our decision to employ the DiNA attention mechanism to promote feature diversity during the learning process and effectively prevent feature collapse.